OpenStudy (anonymous):

Which Property Is NOT Used When Solving 15=2x - ? A) Reflex. Prop. of = B) Add. Prop. of = C)Div. Prop. of = D) Sym. Prop. of =

OpenStudy (across):

What do you mean with\[15=2x-\]It looks like you're missing a term. Either that or the expression is\[15=2x.\]

OpenStudy (anonymous):

Oh, my bad, its 15=2\[x\]-1

OpenStudy (anonymous):

15=2x-1

OpenStudy (across):

The symmetric property is used:\[15=2x-1\iff2x-1=15.\]The additive property is used:\[15=2x-1\implies15+1=2x.\]The division property is used:\[\frac{15+1}{2}=x.\]The reflexive property makes no sense here.